OpenStudy (anonymous):

If f(x)=3x-8 then a) f^-1(y)=? b)f^-1(6)=?

OpenStudy (mertsj):

y=3x-8 Interchange the x and y x=3y-8 Now solve for y x+8=3y y=1/3x+8/3

OpenStudy (mertsj):

Now plug in 6 for y=2+8/3 y=14/3

OpenStudy (anonymous):

Mertsj how can i get x by itself?

OpenStudy (mertsj):

What equation are you referring to?

OpenStudy (anonymous):

1st one

OpenStudy (mertsj):

f(x)=3x-8?

OpenStudy (anonymous):

the y=1/3x+8/3

OpenStudy (mertsj):

Multiply both sides by 3 first of all.

OpenStudy (mertsj):

That gives you: 3y=x+8 Now subtract 8 from both sides and you will have x by itself.

OpenStudy (anonymous):

thanks!

OpenStudy (mertsj):

yw